OpenStudy (anonymous):

simplify to lowest terms. a^2/a+2b-4b^2/a+2b

OpenStudy (raden):

because the denominator was same, just subtract numerators and simplify

OpenStudy (anonymous):

how do ido that

OpenStudy (raden):

like u subtracting 2 fractions 3/4 - 2/4 = (3-2)/4 right ? so, a^2/a+2b - 4b^2/a+2b = (a^2-4b^2)/(a+2b), agree ?

OpenStudy (anonymous):

ok

OpenStudy (raden):

now, can u factorise a^2-4b^2 ?

OpenStudy (anonymous):

how do i factor that

OpenStudy (raden):

a^2-4b^2 = (a+2b)(a-2b) now, apply it to ur fraction. what u get ?

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

Ido not knw how to get the final answer

OpenStudy (raden):

hmmm.. well, look this  hope, this make sense for u